OpenStudy (anonymous):

-7(-81+45)/-6 ~check answer~ P (81+45)= 126 E M D A S -7/126/-6<<<< explain.

OpenStudy (anonymous):

where do i put -7 and -6 in?

OpenStudy (anonymous):

is it \[\frac{-7(-81+45)}{-6}\]

OpenStudy (anonymous):

yeah i got that but hold on so then it would be well if i put it together without seperating it then its a decimal

OpenStudy (anonymous):

yeah i get like -7.4135.....

OpenStudy (anonymous):

Brackets first: \[\frac{-7(-36)}{-6}\] The divide by -6 which we get: \[-7(6)=42\]

OpenStudy (anonymous):

-36/-6=6

OpenStudy (anonymous):

ohhhh okay yeah i was confused about that part

OpenStudy (anonymous):

wait where do you get 36???

OpenStudy (anonymous):

-81+45=-36

OpenStudy (anonymous):

oh nvm